OpenStudy (anonymous):

write the equation of the line containing the given point and parallel the the given line. (5,-9);3x-7y=6. find the equation of the line is y=

OpenStudy (anonymous):

3x - 7y = 6 7y = 3x -6 y = 3x/7 - 6/7 Slope = 3/7. Parallel lines have same slope. So, new line must have same slope = 3/7. Line equation will be: y = 3x/7 + b To find b, you need to plug values of x,y of point 5,-9.